Create the quadratic equation in the form y=ax^2+bx+c using the point (0,2) as one point and the point (9,-8) as the vertex. Enter the a,b, and c values as common fractions in reduced form.
Answer by scott8148(6628)

using the vertex form ___ y = a(x - 9)^2 - 8

substituting ___ 2 = a(0 - 9)^2 - 8 ___ 10 = 81a ___ 10/81 = a

expanding ___ y = (10/81)x^2 - (20/9)x + 2